# Read to Play ([Uncaptioned image] R2-Play): Decision Transformer with Multimodal Game Instruction

### 08 Feb 2024 v2

@@ 1 Introduction
To augment agents’ multitasking and generalization capabilities via multimodal game instructions, leveraging a large, diverse offline dataset for pretraining is vital. This approach is supported by the success of instruction tuning in large language models (Ouyang et al.,, 2022; Brown et al.,, 2020) and large multimodal models (Li et al.,, 2023; Liu et al.,, 2023). The Decision Transformer (DT), an innovative integration of transformers with reinforcement learning (RL), demonstrates superior multitasking capabilities in RL, attributed to its large-scale pretraining  (Chen et al.,, 2021; Lee et al.,, 2022). This advancement establishes DT as a leading approach in the domain of large-scale pretraining for RL. Expanding upon the DT framework, we introduce the Decision Transformer with Game Instruction (DTGI). DTGI is pretrained using Multimodal Game Instructions and a comprehensive offline dataset, designed to simulate the agent’s ability to follow game instructions. Simultaneously, we introduce a novel design called SHyperGenerator, which facilitates knowledge sharing between training and unseen game tasks. Experimental results show that integrating multimodal game instruction into DT leads to a substantial improvement in multitasking and generalization performance. Furthermore, multimodal instruction outperforms textual language and visual trajectory, demonstrating its superior capacity to provide detailed and comprehensive context.
Our contributions are summarized as follows:
@@ 2 Preliminary
In this section, we present the fundamental backgrounds of the Decision Transformer and Hypernetwork.
@@ 3.2 Game Instruction Construction
In the multimodal community, considerable attention is devoted to the integration of instruction tuning within multimodal models  (Xu et al., 2023b, ; Zhang et al., 2023a, ; Li et al.,, 2023; Sun et al., 2023b, ; Sun et al., 2023a, ). These efforts aim to enhance the performance of multimodal models in performing specific visual tasks by incorporating language instructions. Drawing upon the insights derived from these efforts, it is imperative to explore the potential of leveraging multimodal game instructions to augment the capabilities of RL agents, especially in the context of visual-based RL tasks as long-horizon visual tasks. We construct a set of Multimodal Game Instruction (MGI) to apply the benefits of instruction tuning to DT. An illustrative example of a multimodal game instruction is presented in Figure 2.
We introduce a systematic approach for formulating game instructions to explore the incorporation of instruction tuning in RL agents. Initially, we provide ChatGPT  (OpenAI,, 2021) with a thorough overview of the game, encompassing its action space. Following that, we instruct ChatGPT to generate a detailed description for each action, focusing on highlighting key elements. Secondly, we collect gameplay videos featuring expert human players. Subsequently, we downsample and partition these videos into N segments, each segment containing 20 frames. Lastly, we annotate the actions and positions of key elements, represented as [a, b], [c, d]. [a, b] denotes the lower-left point of the element bounding box, while [c, d] denotes the upper-right point of the element’s bounding box.
@@ 4.1 Setup
We select Atari  (Bellemare et al.,, 2013) as our experimental environment for various reasons. Firstly, the visual observations in Atari games consist of pixels, enabling human descriptions of the game screens. Secondly, the action space in Atari is discrete, making it easier to clearly distinguish the agent’s current action. Lastly, Atari offers a diverse range of games with varying distributions, which facilitates the stimulation of the OOD ability of DT. We train our model and the baseline models using a subset of the samples in the DQN-replay dataset, observed by an online agent  (Mnih et al.,, 2015). We evaluate ID ability in 37 training games and 10 unseen games, as reported in Appendix A.1. we report the mean and std of 3 seeds. All raw scores are provided in Appendix A.2. Furthermore, to thoroughly assess the overall performance in multiple games, we standardize the scores. Specifically, we assign the highest positive score as 1 and the lowest negative score as -1. Subsequently, we linearly normalize the scores of all methods to the range of -1 to 1 in each game evaluation.
@@ 4.5 Does the number of training games influence the model’s OOD capabilities?
In this section, we investigate the impact of the quantity of training games on the OOD performance of DTGI. Augmenting the number of training games not only enriches data diversity but also exposes the model to a broader spectrum of instructions, fostering knowledge sharing among them. DTGI undergoes training with three different configurations: 10 games (25%), 20 games (50%), and the entirety of 37 games (100%). The OOD results corresponding to these configurations are elucidated in Table 3.
@@ 4.6 Does DTGI focus on the critical parts of game instruction?
In this section, we investigate the enhancement of Instruction Importance Estimation (see Section 3.3.2) by assigning importance scores to multiple instructions. In Table 1 and 2, we list the results of ID results and OOD results of DTGI and DTGI-a(assign equal scores to each instruction). The instruction set of each game contains 50 instructions. The importance scores for these instructions, visualized in Figure 5, are selectively shown for a subset of games, while comprehensive results for all games are available in Appendix A.4. The analysis specifically highlights the 28th training game, where the score of DTGI is 1.00 and the score of DTGI-a is 0.15.
@@
we observe that Instruction Importance Estimation enhances model performance, as evidenced by the data in Tables 1 and 2. An in-depth examination of the 28th training game demonstrates the correlation between elevated instruction scores and a broader diversity in the trajectory of instruction.
@@ 6 Conclusion
In this paper, we construct a comprehensive multimodal game instruction set, offering extensive context for a variety of game tasks. The experimental findings reveal that integrating multimodal instructions significantly improves performance, outperforming the results obtained with text or trajectory guidance. Additionally, our analysis suggests that broadening the number of games in the dataset enhances the model’s OOD performance more effectively than merely increasing the size of the dataset.
@@ A.3 Experimental Details
Training To ensure fairness and consistency, we maintain uniformity across all baseline models and our model, encompassing both hyper-parameters and the training process. The specific hyper-parameters used for our model and the baseline models are detailed in Table 11. For training purposes, two NVIDIA 4090 graphics cards are utilized in Experiments 4.3, 4.5, and 4.6, while eight NVIDIA 4090 graphics cards are employed in Experiment 4.4.
@@
Batch_size 1000, experiment 4.4; 200, otherwise

### 06 Mar 2024 v3

@@ 1 Introduction
To augment agents’ multitasking and generalization capabilities via multimodal game instructions, leveraging a large, diverse offline dataset for pretraining is vital. This approach is supported by the success of instruction tuning in large language models (Ouyang et al.,, 2022; Brown et al.,, 2020) and large multimodal models (Li et al.,, 2023; Liu et al.,, 2023). The Decision Transformer (DT), an innovative integration of transformers with reinforcement learning (RL), demonstrates superior multitasking capabilities in RL, attributed to its large-scale pretraining  (Chen et al.,, 2021; Lee et al.,, 2022). This advancement establishes DT as a leading approach in the domain of large-scale pretraining for RL. Expanding upon the DT framework, we introduce the Decision Transformer with Game Instruction (DTGI). DTGI is pretrained using Multimodal Game Instructions and a comprehensive offline dataset, designed to simulate the agent’s ability to follow game instructions. Simultaneously, we introduce a novel design called SHyperGenerator, which facilitates knowledge sharing between training and unseen game tasks. Experimental results show that integrating multimodal game instruction into DT leads to a substantial improvement in multitasking and generalization performance. Furthermore, multimodal instruction outperforms textual language and visual trajectory, demonstrating its superior capacity to provide detailed and comprehensive context. Our contributions are summarized as follows:
@@ 2 Preliminary
In this section, we present the backgrounds of the Decision Transformer and Hypernetwork.
@@ 3.2 Game Instruction Construction
In the multimodal community, considerable attention is devoted to the integration of instruction tuning within multimodal models  (Xu et al., 2023b, ; Zhang et al., 2023a, ; Li et al.,, 2023; Sun et al., 2023b, ; Sun et al., 2023a, ). These efforts aim to enhance the performance of multimodal models in performing specific visual tasks by incorporating language instructions. Drawing upon the insights derived from these efforts, it is imperative to explore the potential of leveraging multimodal game instructions to augment the capabilities of RL agents, especially in the context of visual-based RL tasks as long-horizon visual tasks. We construct a set of Multimodal Game Instruction (MGI) to apply the benefits of instruction tuning to DT. An example of multimodal game instruction is presented in Figure 2.
We introduce a systematic approach for formulating game instructions to explore the incorporation of instruction tuning in RL agents. Initially, we provide ChatGPT  (OpenAI,, 2021) with a thorough overview of the game, encompassing its action space. Following that, we instruct ChatGPT to generate a detailed description for each action, focusing on highlighting key elements. Secondly, we collect gameplay videos featuring expert human players. Subsequently, we downsample and partition these videos into N segments, each segment containing 20 frames. Lastly, we annotate the actions and positions of key elements, represented as [a, b], [c, d]. [a, b] denotes the lower-left point of the element bounding box, while [c, d] denotes the upper-right point.
@@ 4.1 Setup
We select Atari  (Bellemare et al.,, 2013) as our experimental environment for various reasons. Firstly, the visual observations in Atari games consist of pixels, enabling human descriptions of the game screens. Secondly, the action space in Atari is discrete, making it easier to clearly distinguish the agent’s current action. Lastly, Atari offers a diverse range of games with varying distributions, which facilitates the stimulation of the OOD ability of DT. We train our model and the baseline models using a subset of the samples in the DQN-replay dataset, observed by an online agent  (Mnih et al.,, 2015). We evaluate ID ability in 37 training games and 10 unseen games, as reported in Appendix A.1. we report the mean and std of 3 seeds. All raw scores are provided in Appendix A.2. Furthermore, to thoroughly assess the overall performance in multiple games, we standardize the scores. Specifically, we assign the highest positive score as 1 and the lowest negative score as -1. Subsequently, we linearly normalize the scores of all methods to the range of -1 to 1.
@@ 4.4 Does the number of training games influence the model’s OOD capabilities?
In this section, we investigate the impact of the quantity of training games on the OOD performance of DTGI. Augmenting the number of training games not only enriches data diversity but also exposes the model to a broader spectrum of instructions, fostering knowledge sharing among them. DTGI undergoes training with three different configurations: 10 games (25%), 20 games (50%), and the entirety of 37 games (100%). The OOD results are elucidated in Table 3.
Table 1: ID results (37 training games)
Table 2: OOD results (10 unseen games)
@@ G DT DTL DTV DTGI-a DTGI
Table 1: ID results (37 training games)
Table 2: OOD results (10 unseen games)
@@ 4.6 Does DTGI focus on the critical parts of game instruction?
In this section, we investigate the enhancement of Instruction Importance Estimation (see Section 3.3.2) by assigning importance scores to multiple instructions. In Table 3 and 3, we list the results of ID results and OOD results of DTGI and DTGI-a(assign equal scores to each instruction). The instruction set of each game contains 50 instructions. The importance scores for these instructions, visualized in Figure 5, are selectively shown for a subset of games, while comprehensive results for all games are available in Appendix A.4. The analysis specifically highlights the 28th training game, where the score of DTGI is 1.00 and the score of DTGI-a is 0.15.
@@
we observe that Instruction Importance Estimation enhances model performance, as evidenced by the data in Tables 3 and 3. An in-depth examination of the 28th training game demonstrates the correlation between elevated instruction scores and a broader diversity in the trajectory of multimodal game instruction.
@@ 6 Conclusion
In this paper, we construct a comprehensive multimodal game instruction set, offering extensive context for a variety of games. The experimental findings reveal that integrating multimodal instructions significantly improves performance, outperforming the results obtained with text or trajectory guidance. Our analysis suggests that broadening the number of games in the dataset enhances the model’s OOD performance more effectively than merely increasing the dataset size.
@@ A.3 Experimental Details
Training To ensure fairness and consistency, we maintain uniformity across all baseline models and our model, encompassing both hyper-parameters and the training process. The specific hyper-parameters used for our model and the baseline models are detailed in Table 11. For training purposes, two NVIDIA 4090 graphics cards are utilized in Experiments 4.3, 4.4, and 4.6, while eight NVIDIA 4090 graphics cards are employed in Experiment 4.5.
@@
Batch_size 1000, experiment 4.5; 200, otherwise